OpenStudy (anonymous):

Hi, I need help in understanding limiting reactants; one problem (an easy one from the looks of it but I don't understand how to do it) is; 1. Sodium metal reacts with oxygen to produce sodium oxide. If 5.00 g of sodium reacted with 5.00 grams of oxygen, how many grams of product is formed? Na(s) + O2(g) -> Na2O(s) Then there's also; How much AgCl product will be produced if 100.00 g of BaCl2 reacted with 100.0 g of AgNO3? AgNO3 + BaCl2 -> AgCl + Ba(NO3)2?

OpenStudy (gebooors):

Balanced equation: 2 Na + O2 -> Na2O First calculate amounts of Na and O2. ( n= m/M) Compare amount of substances to equation: One mole Oxygen needs two moles Na. Is there enough Na? In lower reaction, there should be 2 AgNO3 and 2 AgCl

OpenStudy (anonymous):

I still don't understand how to do it...

OpenStudy (gebooors):

Did you calculate amounts of substances?

OpenStudy (gebooors):

For Na n = 5.0 / 22.99. And for O2 n = 5/32 moles. Is n(Na) more than twice of n(O2)?